DS-Mat regular meeting


Participants: Alex, Bianca, Giorgio, Iza, Maxim, Oscar, Roberto, Susana, Tomasz, Vicente.

Assays status:

connectors for the steering module reported
Ultem 1000 (two weeks measurements): waiting for the report
Sami siltem from Saverio to LNGS (5 kg)

Gd results from LSC. Confirmed the ICPMS results from MSU. Factor 2 worse than Russian purest one, no Po-210 measurement in any case)

Inductors preliminary: 6 Bq/kg.
FEP not doable now.

MC Simulation and R.A. budget updates:

SaG4n rolling. Results to be redone with TENDL to compare with NeuCBOT
Discussion with Vicente Pending.
No progress on the note

Need to evaluate the latest results (Gd and conenctors).

Work on progress on the activation. Possible meeting with Andrea et al. about Ar transportation (low priority), meeting with DEAP collegaues a bout muons UG (high priority)

Surface contamination:
Need to recap the current status (Giorgio/Oscar)
SS sheet ready to be shipped to Legnaro (Cu sheet already at Legnaro)
Meeting with the TPC working tgroup needed to calirfy the SS parts size and welding



There are minutes attached to this event. Show them.
    • 17:00 17:20
      Status and plans 20m
      Speaker: All
    • 17:20 17:30
      Recent assays and new results 10m
      Speaker: All
    • 17:30 17:50
      Status of the R.A. budget 20m
      Speakers: Maxim Gromov (SINP MSU), Vicente Pesudo (CIEMAT)
    • 17:50 17:55
      AOB 5m
      Speaker: All